OpenStudy (anonymous):

14 C has a half-life of 5730 years. If there are 1000 g of 14C initially, how much will be left after 17, 190 years?

OpenStudy (jfraser):

if one half-life is 5730 years, how many half-lives will pass over 17,190?

OpenStudy (anonymous):

The ratio of the left-over stuff (to the original), or n/n0 is (1/2)^(time period/half life). in this case: amount remaining / original amount = (1/2)^(17190 yr/5730 yr) so amount remaining = 1000 grams * (1/2)^(17190 yr/5730 yr)

OpenStudy (anonymous):

I also got 1000 but 1000 isn't on the choices

OpenStudy (aaronq):

hm you can reason these out, but i like to do them mathematically... (These formulas only apply for first order reactions btw) find k, the decay constant, using: \[t _{1/2}=\frac{ \ln2 }{ k }\] the substitute into the exponential decay/growth formula: \[A=A _{0}e ^{-kt} \] t = time elapsed (in terms of half lives) Ao=initial amount (in this case 1000g) A= amount left over after time elapsed (your answer)

OpenStudy (anonymous):

Oh noo! We haven't learned that one... I'm having problems with Chemistry. :( Our teacher had discussed only up to Moles. :((((

OpenStudy (aaronq):

what i wrote is grade 11 math, maybe grade 12.. if you can't do that then apply what the person above me wrote: 1000 grams * (1/2)^(17190 yr/5730 yr)

OpenStudy (anonymous):

I got it already. Thanks.

OpenStudy (aaronq):

good stuff, close the question